OpenStudy (anonymous):

if <1 and <2 are complementary, what are the measures of the angles when m

OpenStudy (anonymous):

Complementary is two angles that sum to 90 degrees. so: 90=(4x+34)+(x+36) 90=4x+34+x+36 90=5x+70 20=5x x=4 So angle a is (4(4)+34) or 50 degrees an angle b is (4+36) or 40 degrees.